# Bounds and Bugs: The Limits of Symmetry Metrics to Detect Partisan Gerrymandering

### v3

@@ Appendix A Proofs of Theorems 1, 3, and 5
Thus, we have so far shown that we can construct election data with minimum MM of
$\frac{1}{2}-V$
For the maximum possible median giving $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$, we would choose $\ell$ districts to have vote share 0, $\frac{n}{2}-\ell$ to have vote share $\frac{1}{2}$ (considered winning districts), and the remaining $\frac{n}{2}$ districts to have vote share $V_{max}$ that results in $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. This $V_{max}$ corresponds to the maximum median. We solve for $V_{max}$ below [^23] This argument glosses over the fact that the median value must be directly in the middle, but by choosing a large number of odd districts and using a limiting argument, we can approach the value stated here.:
$\frac{1}{n}\left(\frac{n}{2}V_{max}+\left(\frac{n}{2}-\ell\right)\frac{1}{2}\right)$ $=V$
$V_{max}=2V-S+\frac{1}{2}$
Of course, for $S<2V-\frac{1}{2}$, this $V_{max}$ is larger than 1, which cannot occur in a district’s vote share. In that case, the largest the median value can be is 1. Thus, for these $(V,S)$, we can construct corresponding election data having
$\frac{1}{2}-V$ $<MM<\min\left\{1-V,V-S+\frac{1}{2}\right\}$
Case 3: $S>\frac{1}{2}$, $\frac{3}{4}\leq V\leq 1$
For $V$ in this range, recall that $S\geq 2V-1$. Thus, we can construct election with half the districts to have vote share 1, $\ell$ to have vote share $\frac{1}{2}$, and the remaining districts to have vote shares resulting in $\frac{1}{n}\sum_{i=1}^{n}V_{i}=V$. This results in seat share $S$ and a maximum median value of 1.
For the minimum possible median giving $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$, we would choose $\ell$ districts to have vote share $\frac{1}{2}$ (corresponding to losing districts), $\frac{n}{2}$ to have vote share $1$ , and the remaining $\frac{n}{2}-\ell$ districts to have vote share $V_{min}$ that results in $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. This $V_{min}$ corresponds to the minimum median. We solve for $V_{min}$ below [^24] Again, we’re glossing over the fact that the median value must be directly in the middle, but by choosing a large number of odd districts and using a limiting argument, we can approach the value stated here.:
$\frac{1}{n}\left(\ell\cdot\frac{1}{2}+\left(\frac{n}{2}-\ell\right)V_{min}+\frac{n}{2}\right)$ $=V$
$\left(\frac{1}{2}-\left(1-S\right)\right)V_{min}$ $=V-\frac{1}{2}-\frac{\ell}{2n}=V-\frac{1}{2}-\frac{1}{2}\left(1-S\right)$
$V_{min}$ $=\frac{2V+S-2}{2S-1}$
Thus, we have shown the following bounds for the Mean-Median Difference:
$\frac{2V+S-2}{2S-1}-V$ $<MM<1-V$

### v4

@@ Appendix A Proofs
has a solution with $p\leq 1$. Thus, we have shown that if $V\leq\frac{3n-2}{4n}$, we can use this construction to obtain election data with MM value of $\frac{1}{2}-V$.
If $V>\frac{3n-2}{4n}$, then the smallest possible median value $p$ would be achieved when $\ell$ districts have vote share $\frac{1}{2}^{-}$, $\frac{n}{2}-1$ districts have vote share 1, and the remaining districts have vote share $p\in{\mathbb{Q}}$ where $p$ satisfies the equation
$\frac{1}{n}\left(\frac{1}{2}\cdot\ell+\left(\frac{n}{2}-\ell+1\right)p+\left(\frac{n}{2}-1\right)\cdot 1\right)=V.$
This results in
$p=\frac{2V+S-2+\frac{2}{n}}{2S-1+\frac{2}{n}}$
and an MM value of $p-V$.
Now suppose that $n$ is odd (so that $n\geq 3$). Then we can choose $\ell$ districts to have vote share $\frac{1}{2}^{-}$ and $\frac{n-1}{2}+1-\ell$ districts to have vote share $\frac{1}{2}^{+}$, giving the smallest possible median value of $\frac{1}{2}$. The remaining districts have vote share $p\in{\mathbb{Q}}$ which results in a statewide vote share of $V$. However this is only possible if the equation
$\frac{\left(\frac{n-1}{2}+1\right)\cdot\frac{1}{2}+p\cdot\frac{n-1}{2}}{n}=V$
has a solution with $p\leq 1$. Thus, we have shown that if $V\leq\frac{3n-1}{4n}$, we can use this construction to obtain election data with MM value of $\frac{1}{2}-V$.
If $V>\frac{3n-1}{4n}$, then the smallest possible median value $p$ would be achieved when $\ell$ districts have vote share $\frac{1}{2}^{-}$, $\frac{n-1}{2}$ districts have vote share 1, and the remaining districts have vote share $p\in{\mathbb{Q}}$ where $p$ satisfies the equation
$\frac{1}{n}\left(\frac{1}{2}\cdot\ell+\left(\frac{n-1}{2}-\ell+1\right)p+\frac{n-1}{2}\cdot 1\right)=V.$
This results in
$p=\frac{2V+S-2+\frac{1}{n}}{2S-1+\frac{1}{n}}$
and a mean-median value of $p-V$.
Case 3: $S>\frac{1}{2}$, $\frac{1}{4}<V\leq 1$, upper bounds for MM
Suppose first that $n$ is even. The maximum possible median is achieved by setting $\ell$ districts to have vote share 0, $\frac{n}{2}-\ell-1$ to have vote share $\frac{1}{2}^{+}$, and the remaining $\frac{n}{2}+1$ districts to have vote share $p\in{\mathbb{Q}}$ where $p$ satisfies the equation
$\frac{1}{n}\left(\left(\frac{n}{2}+1\right)p+\left(\frac{n}{2}-\ell-1\right)\cdot\frac{1}{2}\right)=V.$
This results in
$p=\frac{2V+\frac{1}{2}-S+\frac{1}{n}}{1+\frac{2}{n}}$
and a mean-median value of $p-V$. However, this is only achievable if $p\leq 1$; otherwise the maximum possible median value is 1, resulting in a mean-median value of $1-V$.
Now suppose that $n$ is odd. The maximum possible median is achieved by setting $\ell$ districts to have vote share 0, $\frac{n-1}{2}-\ell$ to have vote share $\frac{1}{2}^{+}$, and the remaining $\frac{n-1}{2}+1$ districts to have vote share $p\in{\mathbb{Q}}$ where $p$ satisfies the equation
$\frac{1}{n}\left(\left(\frac{n-1}{2}+1\right)p+\left(\frac{n-1}{2}-\ell\right)\cdot\frac{1}{2}\right)=V.$
This results in
$p=\frac{2V+\frac{1}{2}-S+\frac{1}{n}}{1+\frac{2}{n}}$
and a mean-median value of $p-V$. However, this is only achievable if $p\leq 1$; otherwise the maximum possible median value is 1, resulting in a mean-median value of $1-V$.
Proof of Theorem 3.
Take the limit as $n\to\infty$ in Theorem 4. ∎